OpenStudy (anonymous):

How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 yards long, and 5 pieces, each 2 yards long, are removed?help me to solve this

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

24 yards intially. You remove 3 yards 3 times and 2 yards 5 times: 24-3*3-2*5=24-9-10=24-19=5

OpenStudy (anonymous):

5 yards left

OpenStudy (gw2011):

24-(3)(3)-(5)(2) 24-9-10=5 yards remain

OpenStudy (anonymous):

the question get wrong its 3 pieces,yards 31/2 and 5 pieces ,yards 21/4 can u please help me to solve this

OpenStudy (gw2011):

24-(3)(3.5)-(5)(2.25) 24-10.5-11.25=2.25 yards remain